OpenStudy (anonymous):

2A/3=8+4A (can't work this out without a calculator) ><

OpenStudy (anonymous):

@A.Avinash_Goutham

OpenStudy (anonymous):

cmon that's too easy.....

sam (.sam.):

\[\frac{2A}{3}=8+4A\] Multiply 3 to both sides 2A=24+12A

OpenStudy (anonymous):

multiply 3 on both sides nd simplify

sam (.sam.):

simplify by subtracting 12A from both sides, then divide

OpenStudy (anonymous):

all this maths is getting to me haha, 13 exams this month...

OpenStudy (anonymous):

good luck :)